OpenStudy (anonymous):

Please help me, Simplify this expression. 4x - 12/X^2 - 9

OpenStudy (anonymous):

factor out a 4 from the numerator = 4(x-3) in the denominator, x^2-9 = (x+3)(x-3) cancel the x-3's you get 4/(x+3)

OpenStudy (anonymous):

Thank you :)